40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed May 13 16:57:28 2009, Dmass, Laser, PMC, Relocking the PMC 
    Reply  Thu May 14 20:57:22 2009, rana, Laser, PMC, Relocking the PMC 
       Reply  Fri May 15 04:23:19 2009, Dmass, Laser, PMC, Relocking the PMC Pmclock.png
Message ID: 125     Entry time: Wed May 13 16:57:28 2009     Reply to this: 127
Author: Dmass 
Type: Laser 
Category: PMC 
Subject: Relocking the PMC 

I have been working at relocking the PMC and have been running into a lot of problems.

 

The table layout when I approached it - the old mode matching lenses were still in place, but the last two the steering mirrors into the PMC after the lenses were gone. I looked at an old elog photo and put the last two mirrors back about at the right place.

 

Then I

  1. Minimized the table power by rotating almost everything into s before the polarizing beam splitter ~100mW of P on the table.
  2. Aligned crudely aligned to the PMC until I got fringes out the other side.
  3. Aligned more by hand, sweeping the cavity with a funciton generator until I started to see what looks like 00, then try to crudely maximize this by hand/eyeball
  4. Then I set up a PDH locking using
  •  A Marconi at 10 dBm and 35 MHz into a splitter
  •  One output of the splitter into the EOM
  •  The other output through a 3 dB attenuator into the L channel of a ZAD-6 Mixer
  •  A Thorlabs PDA10A photodiode (not the best for 1064 - Si) on PMC REFL
  •  The output of the PD into the RF of the Mixer
  • The Output of the Mixer into an SR560 amplifier
  • The output of the SR560 into the DAQ

I was unable to get a good error signal /unable to lock by toggling a 1 Hz pole as I swept through resonance with various DC gains and offsets

 

Things I tried:

  • Changing the PD cable length by 1.4m (about .437*pi)
  • Changing the gain of the SR560 to get in the middle of the dynamic range of my ADC
  • Changing the sign of my feedback to be wrong (desparation?)
  • Checked the alignment onto the photodiode on a scope (tens of mV)
  • Checked the output of the mixer into an HP4395A to make sure the EOM was working - I saw the 35MHz mixed down sideband beat signal - so yes, the eom works.
  • Went to measure the beam waist (maybe my coupling is so bad that even perfectly aligned my error signal would be buried in the noise?) and noticed a serious ghost beam almost coaxial
  • Qualitatively compared the modal content of the beam to an old measurement to see if I had a bimodal PSL output
  • Replaced the first mirror after the mode matching lenses by one with a wedge in a non circular mount.
  • Not using the SR560 and just using an analog passive low pass filter (pole at 1.9 MHz) to kill my mixed down 35 MHz sideband beat
  • Putting an offset on my DAQ input based on what I see as the input when I sweep through 00
  • Using the 35.5 MHz Wenzel frequency doodad (on principle - not as troubleshooting)

 

ELOG V3.1.3-